OpenStudy (anonymous):

Car A has a positive velocity v(t) as it travels along a straight road, where v is a differentiable function of t. The velocity of the car is recorded for several selected values of t over the interval 0<=t<=60 seconds, as shown in the table below t (seconds): 0 10 20 30 40 50 60 v(t) (rate) : 5 14 7 11 12 40 44

OpenStudy (anonymous):

Use the data from the table to approximate the acceleration of Car A at t = 25 seconds Show the computation that leads to the answer please.

OpenStudy (anonymous):

a=du/dt =>a = uf-ui/tf-ti ==>a=30-20/11-7 in m/s^2

OpenStudy (anonymous):

What does "a" represent?

OpenStudy (anonymous):

acceleration

OpenStudy (anonymous):

Ah okay thank you.

OpenStudy (anonymous):

I'm going to go now @surjithayer but I'm still curious as to what you have to say.

OpenStudy (anonymous):

sory my computer failed. doppler is correct.

OpenStudy (anonymous):

Lol okay thanks.